# Towards Fair In-Context Learning with Tabular Foundation Models

### v3

@@ 3.3 In-context Prediction
After performing demonstration selection or transformation using the fairness intervention methods presented previously, we pass them through the tabular foundation model as in-context examples for predicting class labels on the test set. In this paper, we consider three tabular foundation models: TabDPT (Ma et al., 2024), TabICL (Qu et al., 2025), and TabPFN v2 (Hollmann et al., 2025). While all these models use Transformer architectures as their backbones and are pretrained on a large amount of datasets, TabICL and TabPFN use only synthetic data generated from structured causal networks, and TabDPT uses real-world data. The current version of TabPFN can handle a maximum of 10k samples with 500 features, and TabICL can handle up to 500K samples. We randomly subsample in-context examples when the context set size exceeds the maximum number of samples allowed by the foundation model being evaluated.
@@ 4 Experiments
We describe the experimental setup and perform intensive experiments to answer the following research questions: Does correlation remover and group-balanced demonstration selection effectively reduce information about the sensitive attribute and improve the fairness of ICL? What fairness intervention provides a better fairness-accuracy tradeoff? And what foundation tabular (TabPFN vs TabICL) model can provide a better fairness-accuracy tradeoff?
@@ 4.1 Experimental Setup
We experiment on tasks from the recently proposed folktables (Ding et al., 2022), which contains data extracted from the American Community Survey (ACS) Public Use Microdata Sample (PUMS) (Ding et al., 2022). More specifically, we experiment with the following ACS PUMS tasks: ACSIncome, ACSMobility, ACSTravelTime, ACSEmployment, ACSPublicCoverage. These tasks reflect a range of real-world predictive challenges with fairness concerns. A limitation of the ACS PUMS datasets is that they are US-centric; we diversify the experimental setup by including other tasks and datasets. Specifically, we also experiment on the other tabular datasets and tasks including: Diabetes (Gardner et al., 2023), German Credit (Frank, 2010), and CelebA (Liu et al., 2018). More details about each dataset, including the sensitive attributes used for fairness evaluation, the number of samples, and the number of features, can be found in the Appendix A.
@@
In addition to the classic accuracy, we consider three group fairness metrics, i.e., Demographic Parity (DP), Equal Opportunity (EOP), and Equalized Odds (EOD). More details about group fairness metrics can be found in B.1.
@@
For evaluation, we split each dataset into 80%-20%, where the 20% is held out and used to train the sensitive attribute classifier for uncertainty quantification. For the remaining 80% we use 5-fold cross-validation with random shuffle across ten independent runs. This ensures our evaluation is robust and reliable since every data point is used as a test or in-context example across k-fold evaluations. We report the average and standard deviation of fairness and accuracy performance across the k-fold test sets and the ten random independent seeds. As aforementioned, we use TabICL, TabDPT and TabPFN as foundation models with their respective default parameters.
Baselines
We measure and compare the fairness and accuracy performance of ICL under four different in-context selection methods:
Vanilla: the vanilla method performs ICL using randomly selected in-context examples from the training data without any fairness consideration.
Balanced: group-balanced in-context examples selection where context examples are randomly drawn with equal demographic group ratio. The majority group is uniformly downsampled across k-fold evaluations and independent runs.
Correlation Remover: in-context transformation with correlation remover (Feldman et al., 2015) where each non-sensitive feature is transformed to reduce correlation with the sensitive ones. We use the fairlearn toolkit (Bird et al., 2020) implementation and fixed the parameter controlling the fairness-accuracy tradeoff to one, meaning maximal fairness.
Uncertain: using the uncertainty of the sensitive attribute prediction to select in-context examples. We use the Mapie (Taquet et al., 2022) implementation of conformal prediction (Cordier et al., 2023) to estimate the uncertainty of the training data. We fix the coverage parameter , of conformal prediction to and only select as in-context example samples with prediction sets equal to two, i.e., samples with uncertain sensitive attributes. We consider two variants of the method under different model classes used to train the sensitive attribute classifier for uncertainty estimation: a variant that uses the traditional Logistic Regression model (Uncertain+LR) and a variant that uses a foundation model (Uncertain+TabPFN).
@@ 4.2 Results and Discussion
We evaluate several aspects of fairness in ICL prediction with foundational tabular datasets. First, we compare the different baselines considered in terms of fairness and accuracy; For the methods with controllable tradeoff between fairness and accuracy, we vary the hyperparameter controlling the fairness-accuracy tradeoff and compare their Pareto fronts. We then compare the foundation model under the best-performing fairness intervention method. Finally, we provide an ablation study on the impact of the size of the in-context set on fairness and accuracy and discuss the failure case of the correlation removal method.
4.2.1 Baseline Comparison
Figure 1 summarizes the accuracy and fairness outcomes of various baselines across eight datasets, using TabPFN as the foundation model. The group-balanced method yields only marginal improvements in fairness compared to the Vanilla ICL approach, suggesting that representation bias alone does not account for the observed disparities. In other words, even when the in-context sample set is group-balanced, performance disparities across subgroups persist—likely due to other sources of bias such as historical or measurement bias (Mehrabi et al., 2022).
More strikingly, the Correlation Remover method often exacerbates unfairness. For instance, on the ACSIncome dataset, demographic parity disparity increases from to following the application of Correlation Remover. To investigate this failure mode, we assess whether the foundation model can reconstruct the sensitive attribute after Correlation Remover ’s preprocessing step. Specifically, we perform ICL where the sensitive attribute is treated as the prediction target, and measure the model’s test accuracy in reconstructing it.
We find that TabPFN and TabICL can reconstruct the sensitive attribute with up to accuracy (see Table 4 in Appendix), even after Correlation Remover ’s intervention. This suggests that these transformer-based models, pretrained on synthetic data with diverse structural priors, are capable of uncovering and leveraging the transformation patterns introduced by Correlation Remover. Since Correlation Remover modifies each non-sensitive feature based on the sensitive attribute, it unintentionally leaks sensitive information. The foundation model learns to exploit this leakage, ultimately relying on the hidden sensitive signal to predict the target variable—thereby amplifying unfairness.
Further results and discussion are presented in Section 4.5, where we show that applying the feature transformation only to the training set—while keeping the test set unchanged—leads to improved fairness outcomes compared to the Vanilla baseline.
On the other hand, the results also demonstrate that Uncertain methods can significantly improve the fairness of the ICL prediction compared to the different baselines. Table 4 in the Appendix shows that Uncertain methods yield the smallest accuracy of ICL predictions of the sensitive attribute across datasets. This indicates that the selected in-context samples do not encode sufficient information about the sensitive attribute that the foundation model can rely upon for inference, thereby reducing unfairness.
4.2.2 Fairness-Accuracy Tradeoff
In the previous experiment, we fixed the parameters and , which control the tradeoff between fairness and accuracy for the Correlation Remover and Uncertain methods, respectively. To better evaluate this tradeoff, we now vary and over the interval and compare the resulting Pareto fronts of each method.
Using TabPFN as the foundation model, Figures 2(a) and 2(b) display the Pareto fronts on the ACSIncome and ACSMobility datasets, respectively. These figures compare the Vanilla ICL method, Correlation Remover, and the two variants of our uncertainty-based method: Uncertain+LR and Uncertain+TabPFN. Each scatter point corresponds to a different value of or for a given method.
Both variants of the Uncertain method consistently achieve Pareto-dominant points compared to alternatives. In contrast, Correlation Remover can underperform even the Vanilla ICL in terms of fairness, suggesting that it may amplify bias when used with foundation models. Notably, our uncertainty-based method provides a more stable control over the fairness–accuracy tradeoff across values, enabling improved fairness with relatively modest reductions in accuracy.
This accuracy drop is primarily attributed to the smaller in-context set size: as decreases, more samples with high sensitive attribute certainty are excluded from the context, effectively reducing the number of training examples. As discussed further in Section 4.3, the size of the in-context set has a strong influence on model accuracy.
Among the two Uncertain variants, Uncertain+TabPFN —which uses TabPFN for uncertainty estimation—tends to yield better Pareto fronts. This suggests that TabPFN functions as a stronger conformal predictor than logistic regression, allowing for more accurate estimation of sensitive attribute uncertainty and thus more effective in-context selection for fairness.
@@ 4.2.2 Fairness-Accuracy Tradeoff
In this experiment, we extend the comparison to different foundation models by applying the same setup using our best-performing fairness intervention: Uncertain+TabPFN. Figures 3(a) and 3(b) show the fairness–accuracy Pareto fronts on the ACSIncome and ACSMobility datasets, respectively.
Without any fairness intervention (i.e., using the vanilla ICL approach), there is no clear winner among the foundation models in terms of fairness; the results vary across datasets. While one might expect TabDPT —pretrained on real-world data—to encode more real-world biases and thus exhibit poorer fairness performance, this is not consistently observed. In fact, TabDPT often performs comparably to or even better than models pretrained on synthetic data.
Under fairness intervention with Uncertain+TabPFN, all three foundation models exhibit similar fairness peformance, demonstrating that the Uncertain method can consistently control the fairness–accuracy trade-off regardless of the underlying foundation model. However, both TabPFN and TabDPT consistently achieve higher accuracy across datasets. These findings are consistent with previous studies: the authors of TabICL reported strong performance for TabPFN (Qu et al., 2025), while the authors of TabDPT highlighted its competitive performance relative to TabPFN.
We observe similar patterns across additional datasets and results are provided in the Appendix (see Figure 6).
@@ 4.3 Ablation on Impact In-context Sample Size
In all previous experiments, we used the full training set as the in-context example set whenever possible. For example, the current version of TabPFN is limited to handling a maximum of 10,000 samples (Hollmann et al., 2025). To understand the impact of in-context set size, we conduct an ablation study by varying the number of in-context examples across the range [100,300,500,700,1500,2000,2500,3000,4000,5000], while keeping the evaluation setup identical to that described in Section 4.2.1.
Figure 4 shows that accuracy increases substantially with larger in-context sets, particularly in the lower range. However, unfairness shows a slight increase initially and then stabilizes once the in-context set exceeds approximately 700 examples. Across all in-context sizes and datasets, the Correlation Remover method consistently exhibits the highest level of unfairness, reinforcing earlier observations of its bias amplification. In contrast, the Uncertain+TabPFN method consistently achieves the lowest fairness violation, demonstrating robustness to the size of the in-context set. This indicates that even when only a subset of training data can be used, Uncertain+TabPFN remains effective at mitigating bias without overly compromising accuracy.
@@ 4.4 Comparison with LLM-based ICL Methods
Despite LLMs’ in-context learning capabilities, they often lag behind classical machine learning models on large-scale tabular datasets (Hegselmann et al., 2023). To validate the effectiveness of our uncertainty-based demonstration selection method, we compare it against fairness-aware LLM-based fair ICL approaches proposed by Hu et al. (2024) and Bhaila et al. (2024) on the widely studied Adult dataset (Asuncion et al., 2007). Table 1 presents the results. Vanilla TabPFN achieves an accuracy of , significantly outperforming the LLaMA-2-13B baseline, which reaches only . This 10-point gap underscores the performance advantage of specialized models for tabular data, consistent with prior findings that LLMs struggle on such inputs when used naively (Hegselmann et al., 2023).
@@ 4.5 On the Failure of Correlation Remover
In our previous experiments, we observed that applying the correlation removal (CR) to both the training and test data can exacerbate unfairness in ICL predictions. We hypothesized that the foundation model infers the sensitive attribute from the linear transformation applied to each non-sensitive feature, which inadvertently leaks sensitive information and increases unfairness. To validate this hypothesis, we conducted correlation removal prediction of the sensitive attribute after applying fairness interventions. As shown in Table 4, the ICL prediction of the sensitive attribute achieves accuracy following the application of CR. This indicates that the foundation model continues to rely heavily on the sensitive attribute even after correlation removal is performed.
@@
This observation aligns with prior discussions by A¨ıvodji et al. (2021) regarding scenarios where data transformations reduce correlation with sensitive attributes. Besides the classical correlation removal setting (S1), where transformations are applied to both training and test sets, we evaluated the fairness–accuracy trade-off when applying the transformation only to the training set (S2).
@@ 5 Conclusion and future works
In this study, we proposed and analyzed the effectiveness of three preprocessing methods to enhance the fairness of in-context learning (ICL) predictions. Our empirical results, performed on eight fairness benchmarks, posit the uncertainty-based in-context selection method as a strong baseline for improving the fairness of tabular ICL. The key advantages of this method are threefold: (1) it does not require fine-tuning or retraining the foundation model to enforce the desired fairness metrics; (2) it can consistently improve three widely used group fairness metrics; (3) it offers a parameter to control the fairness-accuracy tradeoff. To our knowledge, this is the first work that explores pre-processing fairness intervention on tabular foundation models. We hope this work will trigger more investigations into fair tabular ICL, since in-context learning as a new learning paradigm will be increasingly adopted into decision-making tools. Interesting future research directions include investigating in-processing and post-processing methods and analyzing the effect of distribution shift between in-context and test examples on fairness and accuracy.
Acknowledgement
The authors thank the Digital Research Alliance of Canada and Denvr for computing resources. SEK is supported by CIFAR and NSERC DG (2021-4086) and UA by NSERC DG (2022-04006).

### v4

@@ 3.3 In-Context Prediction
After performing demonstration selection or transformation using the fairness intervention methods presented previously, we pass them through the tabular foundation model as in-context examples for predicting class labels on the test set. In this paper, we consider three tabular foundation models: TabDPT (ma2024tabdpt), TabICL (qu2025tabicl), and TabPFN v2 (hollmann2025accurate). While all these models use Transformer architectures as their backbones and are pretrained on a large amount of datasets, TabICL and TabPFN use only synthetic data generated from structured causal networks, and TabDPT uses real-world data. The current version of TabPFN can handle a maximum of 10k samples with 500 features, and TabICL can handle up to 500K samples. We randomly subsample in-context examples when the context set size exceeds the maximum number of samples allowed by the foundation model being evaluated.
@@ 4 Experiments
We describe the experimental setup and perform intensive experiments to answer the following research questions:
(R1) Does group-balanced demonstration selection and correlation remover effectively reduce information about the sensitive attribute and improve the fairness of ICL?
(R2) What fairness intervention provides a better fairness-utility tradeoff?
(R3) What foundation tabular model can provide a better fairness-utility tradeoff?
@@ 4.1 Experimental Setup
We experiment on tasks from folktables (ding_retiring_2022), which contains data extracted from the American Community Survey (ACS) Public Use Microdata Sample (PUMS). More specifically, we experiment with the following ACS PUMS tasks: ACSIncome, ACSMobility, ACSTravelTime, ACSEmployment, ACSPublicCoverage. These tasks reflect a range of real-world predictive challenges with fairness concerns. We use the data of the year 2018 from the state of Alabama (AL), which is one of the states with the largest fairness violation (ding_retiring_2022). A limitation of the ACS PUMS datasets is that they are US-centric; we diversify the experimental setup by including other tasks and datasets. Specifically, we also experiment on the other tabular datasets and tasks including: Diabetes (gardner2023tableshift), German Credit (frank2010uci), and CelebA (liu2018large). More details about each dataset, including the sensitive attributes used, the number of samples, and the number of features, can be found in the Appendix A.
@@
We measure the utility of the model using accuracy, and measure fairness using three group fairness metrics, i.e., Demographic Parity (DP), Equal Opportunity (EOP), and Equalized Odds (EOD). More details about group fairness metrics can be found in B.1.
Baselines
We evaluate ICL under four in-context selection strategies, comparing both fairness and accuracy:
Vanilla: randomly selects in-context examples from the training set without fairness considerations.
Group Balanced: samples in-context examples to maintain equal group ratios, uniformly downsampling the majority group across runs.
Correlation Remover (CR): applies correlation-removal (feldman2015certifying) to reduce dependence between sensitive and non-sensitive features in both in-context and test examples. We use the fairlearn implementation (bird2020fairlearn) with varying values of , where enforces maximal fairness.
Uncertain: selects in-context examples with high uncertainty in sensitive attribute prediction. We estimate uncertainty using conformal prediction (Cordier_Flexible_and_Systematic_2023) as implemented in Mapie (taquet2022mapie), varying the coverage parameter to control the fairness–utility tradeoff. We study two variants of sensitive attribute classifier: one with a Logistic Regression classifier (Uncertain+LR) and one with a foundation model (Uncertain+TabPFN).
@@
For evaluation, we held out 20% of the data to train the sensitive attribute classifier used in the uncertainty-based baseline. The remaining data were split into 80% for training and 20% for testing. We applied the fairness interventions to the training set to derive in-context samples, and reported fairness and accuracy on the test set, averaged over three random seeds. This setup provides robustness, as each data point can appear as a test sample or an in-context sample across seeds. As noted above, we use TabICL, TabDPT, and TabPFN with their default parameters.
@@ 4.2 Results and Discussion
We evaluate fairness in ICL predictions with tabular foundation models along several dimensions. First, we compare the baseline methods in terms of fairness and accuracy. For interventions with a tunable fairness–utility tradeoff, we vary the corresponding hyperparameter and compare their Pareto fronts. While we use accuracy as the main utility metric, we verify in the Appendix that results are consistent when using ROC AUC. Next, we compare foundation models under the best-performing intervention. Finally, we conduct an ablation on the effect of in-context set size and analyze the failure case of the correlation-removal method.
4.2.1 Fairness-Utility Tradeoff
Figure 1 shows the fairness–utility Pareto fronts on ACSIncome, ACSMobility, and ACSTravelTime datasets using TabPFN as the foundation model. We compare the Vanilla baseline, Group Balanced, Correlation Remover, and the two variants of our uncertainty-based method (Uncertain+LR and Uncertain+TabPFN). For Correlation Remover and Uncertain, we vary the tradeoff parameters and over , with each shaded point corresponding to a parameter setting. The main findings, consistent across datasets and models, are as follows:
(R1) The Group Balanced method yields only marginal fairness gains over Vanilla ICL, with similar utility. This suggests that representation bias alone does not explain observed disparities; even with group-balanced contexts, subgroup performance gaps remain, likely due to other sources of bias such as historical, algorithmic, or measurement bias (mehrabi_survey_2022).
(R2) In contrast, both variants of Uncertain consistently achieve Pareto-dominant points. They offer stable control over the fairness–accuracy tradeoff across values, enabling significant fairness improvements with modest accuracy loss. The reduced accuracy stems mainly from smaller in-context sets, since excluding low-uncertainty samples decreases the size of the in-context set, which can reduce accuracy as we discuss in Section 4.3).
(R2, Uncertainty variants) Among the two uncertainty-based methods with different model classes, Uncertain+TabPFN typically yields stronger Pareto fronts, indicating that TabPFN provides more reliable conformal predictions than logistic regression, leading to better fairness-aware sample selection.
4.2.2 Accuracy of sensitive attribute reconstruction
| Dataset | ICL Method | TabPFN | TabICL | ||
| Accuracy | F1 Score | Accuracy | F1 Score | ||
| ACSIncome | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSTravelTime | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSPublicCoverage | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSEmployment | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| ACSMobility | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| German Credit | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| Diabetes | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
| CelebA | Vanilla | ||||
| Group Balanced | |||||
| Correlation R. | |||||
| Uncertain+LR | |||||
| Uncertain+TabPFN | |||||
We assess whether foundation models can reconstruct the sensitive attribute after different fairness interventions. Specifically, we perform ICL to predict the sensitive attribute and measure test accuracy. Effective debiasing should lower this accuracy, ideally approaching random guessing ().
Table 1 shows that the Group Balanced intervention only marginally reduces reconstruction accuracy for TabPFN and TabICL on ACS datasets. This limited effect is expected, since ACS is nearly group-balanced for gender (ding_retiring_2022), e.g., for the state of Alabama, the gender distribution is females, males. In contrast, the Diabetes dataset is highly imbalanced by race, and group balancing reduces reconstruction accuracy from for the vanilla baseline to .
For Correlation Remover, we observe the opposite effect: reconstruction accuracy rises to nearly . We identify this as the source of bias amplification. Transformer-based foundation models, pretrained on synthetic data with diverse structural priors, appear to detect and exploit the transformations applied by Correlation Remover. Because each non-sensitive feature is modified using the sensitive attribute, sensitive information leaks into the transformed features. The models leverage this leakage, relying on hidden sensitive signals to predict the target variable and thereby amplifying unfairness. We provide further analysis in Section 4.5, showing that applying correlation removal only to the training set—while keeping the test set unchanged—improves fairness compared to the Vanilla baseline.
Finally, the Uncertain methods yield the lowest reconstruction accuracy across datasets. This suggests that the selected in-context samples carry little information about the sensitive attribute, limiting the model’s ability to exploit it and thereby improving fairness.
@@ 4.2.2 Accuracy of sensitive attribute reconstruction
In this experiment, we extend the comparison to different foundation models by applying the same setup using our best-performing fairness intervention: Uncertain+TabPFN. Figures 2(a) and 2(b) show the fairness–accuracy Pareto fronts on the ACSIncome and ACSMobility datasets, respectively.
Without any fairness intervention (i.e., using the vanilla ICL approach), there is no clear winner among the foundation models in terms of fairness; the results vary across datasets. While one might expect TabDPT —pretrained on real-world data—to encode more real-world biases and thus exhibit poorer fairness performance, this is not consistently observed. In fact, TabDPT often performs comparably to or better than foundation models pretrained on synthetic data (R3).
Under fairness intervention with Uncertain+TabPFN, all three foundation models exhibit similar fairness performance, demonstrating that the Uncertain method can consistently control the fairness–accuracy trade-off regardless of the underlying foundation model. However, both TabPFN and TabDPT consistently achieve higher accuracy across datasets (R3). These findings are consistent with previous studies: the benchmark erickson_tabarena_2025 reported strong performance for TabPFN (qu2025tabicl), while the authors of TabDPT highlighted its competitive performance relative to TabPFN. We observe similar patterns across additional datasets, and results are provided in the Appendix (see Figure 5).
@@ 4.3 Ablation on Impact In-context Sample Size
In all previous experiments, we used the full training set as the in-context example set whenever possible. For example, the current version of TabPFN is limited to handling a maximum of 10,000 samples (hollmann2025accurate). To understand the impact of in-context set size, we conduct an ablation study by varying the number of in-context examples across the range [100,300,500,700,1500,2000,2500,3000,4000,5000], while keeping the evaluation setup identical as described above.
Figure 3 shows that accuracy increases substantially with larger in-context sets, particularly in the lower range. However, unfairness shows a slight increase initially and then stabilizes once the in-context set exceeds approximately 700 examples. Across all in-context sizes and datasets, the Correlation Remover method consistently exhibits the highest level of unfairness, reinforcing earlier observations of its bias amplification. In contrast, the Uncertain+TabPFN method consistently achieves the lowest fairness violation, demonstrating robustness to the size of the in-context set. This indicates that even when only a subset of training data can be used, Uncertain+TabPFN remains effective at mitigating bias without overly compromising accuracy. We observe a similar trend when using TabICL model, and results are provided in Appendix 13.
In Appendix C, we perform a computational analysis comparing running and inference time of different fainress interventions and foundation models. This analysis shows that fairness interventions themselves are computationally inexpensive. The runtime is dominated by the choice of the tabular foundation model. TabPFN is the most efficient, benefiting from its compact architecture. The uncertainty-based variant with LR is fastest, while using it with TabPFN itself for uncertainty estimation adds moderate overhead..
@@ 4.4 Comparison with LLM-based ICL Methods
| Method | Foundation model | Accuracy | ||
| Vanilla | TabPFN | |||
| Vanilla | LLaMA-2-13B | |||
| hu2024strategic | GPT-3.5-turbo | |||
| bhaila2024fair | LLaMA-2-13B | |||
| Uncertain+TabPFN | TabPFN |
Despite LLMs’ in-context learning capabilities, they often lag behind classical machine learning models on large-scale tabular datasets (hegselmann2023tabllm). To validate the effectiveness of our uncertainty-based demonstration selection method, we compare it against fairness-aware LLM-based fair ICL approaches proposed by hu2024strategic and bhaila2024fair on the widely studied Adult dataset (asuncion2007uci). Table 2 presents the results. Vanilla TabPFN achieves an accuracy of , significantly outperforming the LLaMA-2-13B baseline, which reaches only . This 10-point gap underscores the performance advantage of specialized models for tabular data, consistent with prior findings that LLMs struggle on such inputs when used naively (hegselmann2023tabllm).
@@ 4.5 On the Failure of Correlation Remover
In our previous experiments, we observed that applying the CR to both the training and test data can exacerbate unfairness in ICL predictions. We hypothesized that the foundation model infers the sensitive attribute from the linear transformation applied to each non-sensitive feature, which inadvertently leaks sensitive information and increases unfairness. To validate this hypothesis, we conducted correlation removal prediction of the sensitive attribute after applying fairness interventions. As shown in Table 1, the ICL prediction of the sensitive attribute achieves accuracy following the application of CR. This indicates that the foundation model continues to rely heavily on the sensitive attribute even after correlation removal is performed.
@@
This observation aligns with prior discussions by aivodji2021local regarding scenarios where data transformations reduce correlation with sensitive attributes. Besides the classical correlation removal setting (S1), where transformations are applied to both training and test sets, we evaluated the fairness–accuracy trade-off when applying the transformation only to the training set (S2).
@@ 5 Conclusion and future works
In this study, we proposed and analyzed the effectiveness of three preprocessing methods to enhance the fairness of in-context learning (ICL) predictions. Our empirical results, performed on eight fairness benchmarks, posit the uncertainty-based in-context selection method as a strong baseline for improving the fairness of tabular ICL. The key advantages of this method are threefold: (1) it does not require fine-tuning or retraining the foundation model to enforce the desired fairness metrics; (2) it can consistently improve three widely used group fairness metrics; (3) it offers a parameter to control the fairness-utility tradeoff. To our knowledge, this is the first work that explores pre-processing fairness intervention on tabular foundation models. We hope this work will trigger more investigations into fair tabular ICL, since in-context learning as a new learning paradigm will be increasingly adopted into decision-making tools. Interesting future research directions include investigating in-processing and post-processing methods and analyzing the effect of distribution shift between in-context and test examples on fairness and accuracy.